# SynCL: A Synergistic Training Strategy with Instance-Aware Contrastive Learning for End-to-End Multi-Camera 3D Tracking

### v2

@@ Abstract
While existing query-based 3D end-to-end visual trackers integrate detection and tracking via the tracking-by-attention paradigm, these two chicken-and-egg tasks encounter optimization difficulties when sharing the same parameters. Our findings reveal that these difficulties arise due to two inherent constraints on the self-attention mechanism, i.e., over-deduplication for object queries and self-centric attention for track queries. In contrast, removing self-attention mechanism not only minimally impacts regression predictions of the tracker, but also tends to generate more latent candidate boxes. Based on these analyses, we present SynCL, a novel plug-and-play synergistic training strategy designed to co-facilitate multi-task learning for detection and tracking. Specifically, we propose a Task-specific Hybrid Matching module for a weight-shared cross-attention-based decoder that matches the targets of track queries with multiple object queries to exploit promising candidates overlooked by the self-attention mechanism. To dynamically select optimal candidates for the one-to-many matching, we also design a Cost-based Query Filtering module controlled by model training status. Moreover, we introduce Instance-aware Contrastive Learning to break through the barrier of self-centric attention for track queries, effectively bridging the gap between detection and tracking. Extensive experiments demonstrate that SynCL consistently delivers improvements when integrated with the training stage of various query-based 3D MOT trackers. Without additional inference costs, SynCL improves the state-of-the-art PF-Track method by $+3.9\%$ AMOTA and $+2.0\%$ NDS on the nuScenes dataset.
@@ 1 Introduction
Figure 1: Comparisons between the tracking-by-attention paradigm and our proposed plug-and-play training strategy, SynCL. SynCL consists of Task-specific Hybrid Matching, Instance-aware Contrastive Learning powered by a Cost-based Query Filtering moudle.
Figure 2: We compare results from the standard decoder (Fig.2a) with those from the decoder without self-attention (Fig.2b) in inference stage. The self-attention mechanism exhibits over-deduplication for object queries and self-centric attention for track queries.
@@
The perception system is an indispensable component in autonomous driving. Within the system, accurate 3D multi-object tracking (MOT) provides reliable observations for planning. Camera-based tracking algorithms have raised significant attention due to their cost-effectiveness Hu et al. (2023a). Traditional trackers following the tracking-by-detection paradigm often suffer from tedious and unstable post-processing in new scenarios Wang et al. (2024). To achieve cross-domain adaptability, the prevailing methods Zhang et al. (2022b); Pang et al. (2023); Ding et al. (2024) adopt the tracking-by-attention paradigm, which integrates single-frame detection and inter-temporal tracking by encoding new-born and persistent targets as object queries and track queries respectively (Fig.1a). However, the optimization of detection and tracking, two interdependent yet distinctly characterized tasks, remains an open question due to the sharing of the same model parameters within the end-to-end training pipeline Zhang et al. (2023); Yu et al. (2023). Previous works attribute the optimization difficulties to conflicts in representation and gradient flows across distinct queries. DQTrack Li et al. (2023) attempts to train detection and learning-based association separately, but this two-stage training pipeline fails to enable synergistic representation learning. ADATrack Ding et al. (2024) introduces an additional attention-based association module to strengthen inter-query connections, yet consequently increases model complexity. OneTrack Wang et al. (2024) groups object queries and blocks conflicting gradient flows, but its design is confined to a specific detector, thus lacking generalizability. A question is: Is there a universal training strategy that can effectively address the optimization difficulties of detection and tracking under shared model parameters through a joint training paradigm without affecting inference speed?
@@
To seek the answer, we first conduct a comprehensive analysis of attention mechanisms in the tracking-by-attention paradigm. We find that the removal of the self-attention mechanism from the standard decoder does not significantly affect the regression predictions for both object and track queries (Fig.2b) compared to results from the standard decoder (Fig.2a). Interestingly, the self-attention mechanism tends to de-duplicate the predictions from object queries that are similar to those from track queries. Although this de-duplication plays the similar role of Non-Maximum Suppression (NMS) to some extent, it inadvertently results in information loss, potentially eliminating some high-quality candidates (e.g., ID 2). Moreover, track queries are constrained in self-centric attention, lacking precise inter-temporal localization capabilities (e.g., ID 5, ID 8, etc.). Despite the above issues, directly replacing the self-attention mechanism with NMS will lead to the collapse of joint training. In another way, optimization difficulties may be mitigated by a weight-shared cross-attention-based parallel decoder, following Hu et al. (2023b). However, simultaneously addressing the over-deduplication of object queries and the self-centric issues of track queries within a parallel decoder remains non-trivial.
@@
To thoroughly tackle these challenges, we propose a novel synergistic training strategy for the two tasks based on the tracking-by-attention paradigm, which we term SynCL, presented in Fig.1b. Concretely, we propose a Task-specific Hybrid Matching module that additionally employs one-to-many assignment for object queries to uncover and preserve promising candidates overlooked by the self-attention mechanism in the parallel decoder. With more diverse and refined candidate representations from one-to-many label supervision, predictions from object queries are robust to over-deduplication. Additionally, we design a Cost-based Query Filtering module based on the Gaussian Mixture Model (GMM) that dynamically selects reliable object queries for the one-to-many assignment according to the prediction quality and optimization state of the model. Moreover, to get rid of constraints from self-centric attention on track queries and establish cross-task connections, we introduce Instance-aware Contrastive Learning that aligns joint object and track queries corresponding to identical ground truth targets while separating irrelevant queries in the latent space and thus provide synergistic learning with richer and quality features.
In summary, our main contributions are as follows:
  • We reveal that the optimization difficulties of joint detection and tracking training lie in the discriminative effect of the self-attention mechanism across different queries.
@@
  • We propose SynCL, a synergistic training strategy that can be compatible with any tracking-by-attention paradigm tracker to address multi-task learning challenges. SynCL implements dynamic filtering-based hybrid matching and instance-aware contrast learning to improve the performance of both detection and tracking.
  • We empirically show that SynCL brings remarkable improvements over strong tracking-by-attention baselines with acceptable extra training cost and performs favorably against recent state-of-the-art methods. Code and raw results will be publicly available upon acceptance.
@@ 2 Related Work
2.1 Tracking-by-detection
Traditional tracking methods in 2D or 3D Weng et al. (2020a); Pang et al. (2022); Zhang et al. (2022c); Wojke et al. (2017) follow the Tracking-by-detection paradigm, which decouples detection and tracking, regarding tracking as a post-processing step. Early works associate targets with existing trajectories by handcrafted features such as appearance similarity Hu et al. (2023a), geometric distance Yin et al. (2021); Bai et al. (2022); Zhou et al. (2019). Tracking-by-detection paradigm achieves further advancements when combined with query-based detectors. Specifically, learning modules such as GNN Weng et al. (2020c, 2021) and Transformers Zeng et al. (2021); Li et al. (2023) are employed to merge multifaceted information. However, detection and tracking are complementary yet non-independent tasks. This partitioned paradigm design does not effectively optimize both tasks simultaneously.
2.2 Multi-task learning in MOT
Building upon DETR Carion et al. (2020) of encoding detection targets as object queries, MOTR Zeng et al. (2022) utilizes the track queries, enriched with prior features of object queries, to propagate across frames, enabling an end-to-end tracking framework. MUTR3D Zhang et al. (2022b) and PF-Track Pang et al. (2023) extend the end-to-end framework to multi-camera 3D MOT. However, sharing the same model parameters for detection and tracking within an end-to-end framework introduces optimization difficulties in multi-task learning. Subsequent works Zhang et al. (2023); Yu et al. (2023); Li et al. (2023) attempt to address these difficulties. STAR-Track Doll et al. (2024) introduces a Latent Motion Model (LMM) to account for changes in query appearance features across frames and ADA-Track Ding et al. (2024) integrates the appearance and geometric clues of the track queries to refine object queries. Both of them increase inference speed. OneTrack Wang et al. (2024) utilizes attention mask to block conflicting gradient flows, but its design is confined to StreamPETR Wang et al. (2023). Despite these efforts, a generic strategy for addressing optimization difficulties in joint detection and tracking training remains an open problem.
@@ 3.2 Preliminaries
SynCL is designed to be integrated into any tracking-by-attention paradigm multi-view 3D tracker. These trackers typically consist of a convolutional network-based feature encoder (e.g., ResNet He et al. (2016), VoVNet Lee et al. (2019)), a transformer-based decoder, and a predictor head for object classes and boxes. The goal of 3D MOT is to generate detection boxes of the instances in the perceptual scene with consistent IDs. Given $N$ images $\mathbf{I}_{t}=\{\mathbf{I}_{t}^{i},i=1,2,\dots,N\}$ captured by surrounding cameras at timestamp $t$, the encoder first extracts image features $\mathbf{F}_{t}=\{\mathbf{F}_{t}^{i},i=1,2,\dots,N\}$ with 3D point position embeddings Liu et al. (2022); Wang et al. (2021). The key point in realizing end-to-end tracking is to allow the prior latent encoding from the previous frame, track queries $\mathbf{Q}^{trk}_{t-1}$, to be updated iteratively by the decoder interacting with the current frame image features. To detect new-born targets, object queries $\mathbf{Q}^{obj}$ with a fixed number of $\mathbf{N}_{obj}$ are merged together and fed into the decoder:
@@
However, matching object queries to new-born objects contradicts with detection pre-training stage, substantially weakening the supervision signals for object queries. Consequently, when track queries are initialized with object queries from the previous frame, these low-quality priors inevitably exacerbate the optimization difficulties in inter-temporal tracking. We found that the suboptimal assignment stems from the discriminative nature of the self-attention mechanism, preventing synergistic training for both queries.
Figure 3: Analysis of self-attention heatmap in the standard decoder.
Figure 4: Overview of SynCL. SynCL is based on tracking-by-attention paradigm trackers, with two weight-share parallel decoders: a S-decoder (standard decoder) and a C-decoder (devoid of self-attention layers). In C-decoder, hybrid matching with one-to-many and one-to-one assignment is applied for object queries and track queries, respectively. Besides, a cost-based filtering module is designed to dynamically select reliable object queries for the one-to-many assignment. With identical ground-truth matching, contrastive learning further unify the representations between object and track queries, co-facilitating multi-task learning for detection and tracking.
@@ 3.3 Task-specific Hybrid Matching
3.3.1 Parallel Decoders
We design two parallel decoders: a standard decoder (S-decoder) and a cross-attention-based decoder (C-decoder). A block in the standard decoder consists of a self-attention layer, a cross-attention layer and a feed-forward network (FFN). The embeddings of both object and track queries are concatenated and fed into the two decoders, denoted as $e^{l}_{s}$ and $e^{l}_{c}$, where $e^{l}=e^{l}_{obj}\cup e^{l}_{trk}$. The superscript $l=\{0,1,...,5\}$ indicates the index of the decoder block and we omit the notation of the frame index $t$ for simplicity. The initial inputs to both decoders are identical, i.e., $e^{0}_{s}=e^{0}_{c}$. In the $l_{th}$ block of the S-decoder, the state of both queries is first updated through the self-attention layer bidirectionally:
$\hat{e}^{l}_{s}=e^{l}_{s}+{\rm SA}(Q=K=V=e^{l}_{s})$ (3)
where $\rm SA(\cdot)$ indicates the self-attention layer. Next, the two types of queries are interacted with multi-view image features, where object queries generate potential candidates through the positional priors, and track queries locate persistent targets using discriminative representation prior:
$e^{l+1}_{s}={\rm FFN}(\hat{e}^{l}_{s}+{\rm CA}(Q=\hat{e}^{l}_{s},K=V=\mathbf{F}))$ (4)
where $\rm CA(\cdot)$ indicates the cross-attention layer. The C-decoder retains the cross-attention layers and FFNs, sharing parameters with the S-decoder. In this setup, the cross-attention layers take image features as the sole source set (key and value), dynamically aggregating and updating both queries in the stacked decoder blocks:
$e^{l+1}_{c}={\rm FFN}(e^{l}_{c}+{\rm CA}(Q=e^{l}_{c},K=V=\mathbf{F}))$ (5)
Due to the absence of mutual interaction between object and track queries, it is challenging for C-decoder to remove duplicate candidates with proximal or occluded targets. Thus, we design distinct label assignment tailored to the characteristics of the two parallel decoders and the specific requirements of their respective tasks.
3.3.2 Hybrid Matching
End-to-end tracking-by-attention paradigm trackers rely on one-to-one assignment for both object and track queries. In the S-decoder, assignment of object queries $\mathbf{\sigma}^{obj}_{s}$ can be obtained by performing the Hungarian algorithm between predictions and new-born targets, while assignment of track queries $\mathbf{\sigma}^{trk}_{s}$ is consistent with the assigned IDs:
@@ 3.3.2 Hybrid Matching
$\mathbf{\sigma}^{trk}_{s}$ =$=\left\{\sigma:\mathrm{ID}_{\sigma(i)}^{trk}=\mathrm{\hat{ID}}_{i},\mathrm{\hat{ID}}\in g_{pst},i=1,2,...,N\right\}$ (7)
Where $g_{new}$ and $g_{pst}$ are $M$ new-born and $N$ persistent ground truths (GTs). $\mathcal{C}_{m}$ is the matching cost between predictions and GTs. $\sigma(\cdot)$ is the optimal permutation of $M$ or $N$ indices, where $k_{th}$ ground truth target is assigned to $\sigma(k)_{th}$ prediction. Each query corresponds to a ground-truth target and vice versa. The track queries corresponding to disappearing targets are not assigned labels. Following Wang et al. (2021); Liu et al. (2023), we utilize a combination of classification score and the 3D box as the matching cost metric:
$C_{m}={\rm FocalLoss}(p,\hat{p})+{\rm L1}(b,\hat{b})$ (8)
@@
Where $K$ is the maximum assignment constraint per GT for object queries. In order to obtain more positive sample pairs for contrastive learning, we do not match object queries with all ground-truth targets. In Sec.4.3, we explore the impact of different label assignment variants in detail.
3.4 Cost-based Query Filtering
As the training iterations progress, the parallel decoder based on cross-attention mechanism gradually improves the quality of its predictions. To more effectively explore high-quality candidate samples, we propose a cost-based filtering module with a dynamic threshold, which intelligently constructs a progressive learning strategy ranging from simple to complex according to the model’s optimization status. Specifically, we classify the query-to-ground-truth cost distribution in the C-decoder into reliable and unreliable states, and apply a Gaussian Mixture Model (GMM) to unsupervisedly cluster matching sample pairs. The GMM is defined as:
$\mu,\Sigma=\arg\max\sum_{i=1}^{N_{c}}\pi_{i}\mathcal{N}(z|\mu_{i},\Sigma_{i})$ (10)
where $\mathcal{N}(z|\mu_{i},\Sigma_{i})$ is the Gaussian probability density function with mean $\mu_{i}$ and covariance $\Sigma_{i}$, and $z$ is the random variable of the cost value. We have $N_{c}=2$ Gaussian components and $\pi_{i}$ is the weight for $i$-th component, $\sum_{i=1}^{N_{c}}\pi_{i}=1$. We further set the clustering center of reliable ones $\mu_{reliable}$ as the threshold $\tau$. We filter out unreliable matching pairs whose cost value exceeds the dynamic threshold $\tau$.

### v3

@@ Abstract
While existing query-based 3D end-to-end visual trackers integrate detection and tracking via the tracking-by-attention paradigm, these two chicken-and-egg tasks encounter optimization difficulties when sharing the same parameters. Our findings reveal that these difficulties arise due to two inherent constraints on the self-attention mechanism, i.e., over-deduplication for object queries and self-centric attention for track queries. In contrast, removing the self-attention mechanism not only minimally impacts regression predictions of the tracker, but also tends to generate more latent candidate boxes. Based on these analyses, we present SynCL, a novel plug-and-play synergistic training strategy designed to co-facilitate multi-task learning for detection and tracking. Specifically, we propose a Task-specific Hybrid Matching module for a weight-shared cross-attention-based decoder that matches the targets of track queries with multiple object queries to exploit promising candidates overlooked by the self-attention mechanism. To flexibly select optimal candidates for the one-to-many matching, we also design a Dynamic Query Filtering module controlled by model training status. Moreover, we introduce Instance-aware Contrastive Learning to break through the barrier of self-centric attention for track queries, effectively bridging the gap between detection and tracking. Without additional inference costs, SynCL consistently delivers improvements in various benchmarks and achieves state-of-the-art performance with 58.9% AMOTA on the nuScenes dataset. Code and raw results will be publicly available.
@@ 1 Introduction
The perception system is an indispensable component in autonomous driving. Within the system, accurate 3D multi-object tracking (MOT) provides reliable observations for planning. Camera-based tracking algorithms have raised significant attention due to their cost-effectiveness [1]. Traditional trackers following the tracking-by-detection paradigm often suffer from tedious and unstable post-processing in new scenarios [2, 3]. To achieve cross-domain adaptability, the prevailing methods [4, 5, 6, 7, 8] adopt the tracking-by-attention paradigm, which integrates single-frame detection and inter-temporal tracking by encoding new-born and persistent targets as object queries and track queries respectively (Fig. 1a). However, the optimization of detection and tracking, two interdependent yet distinctly characterized tasks, remains an open question due to the sharing of the same model parameters within the end-to-end training pipeline [9, 10]. Previous works attribute the optimization difficulties to conflicts in representation [6] and gradient flows across distinct queries [8]. DQTrack [11] attempts to train detection and learning-based association separately, but this two-stage training pipeline fails to enable synergistic representation learning. ADATrack [6] introduces an additional attention-based association module to strengthen inter-query connections, yet consequently increases model complexity. OneTrack [8] groups object queries and blocks conflicting gradient flows, but its design is confined to a specific detector, thus lacking generalizability. A question is: Is there a universal training strategy that can effectively address the optimization difficulties of detection and tracking under shared model parameters through a joint training paradigm without affecting inference speed?
Figure 1: Comparisons between the tracking-by-attention paradigm and our proposed plug-and-play training strategy, SynCL. SynCL consists of Task-specific Hybrid Matching, Instance-aware Contrastive Learning powered by a Dynamic Query Filtering moudle.
Figure 2: We compare results from the standard decoder in Fig. 2b with those from the decoder without self-attention in Fig. 2c in inference stage. The self-attention mechanism exhibits over-deduplication for object queries and self-centric attention for track queries. Results are from the model trained without utilizing SynCL.
To seek the answer, we first conduct a comprehensive analysis of attention mechanisms in the tracking-by-attention paradigm. We find that the self-attention mechanism exhibits two inconspicuous characteristics for object queries and track queries, respectively:
  • Over-deduplication for object queries: Referring to the purple dashed box in Fig. 2a, the self-attention mechanism in the decoder assigns high attention scores from object queries to similar track queries. Compared to the prediction results before and after removing the self-attention mechanism (see purple circles in Fig. 2b and Fig. 2c), we observe that these high attention scores effectively suppress duplicate predictions (e.g., ID 0, ID 1 and ID 4). Although this de-duplication plays a similar role of Non-Maximum Suppression (NMS) to some extent, it inadvertently results in information loss, potentially eliminating some high-quality candidates (e.g., ID 2).
  • Self-centric attention for track queries: Referring to the green dashed box in Fig. 2a, track queries exhibit high attention scores towards themselves, lacking interaction with object queries. As illustrated in the green circles of Fig. 2b and Fig. 2c, the imprecise predictions of track queries (e.g., ID 6, ID 8, etc.) change little before and after the removal of the self-attention mechanism.
@@
Despite the above issues, directly replacing the self-attention mechanism with NMS will lead to the collapse of joint detection and tracking training. In another way, optimization difficulties may be mitigated by a weight-shared cross-attention-based parallel decoder, following [12]. However, simultaneously addressing the over-deduplication of object queries and the self-centric issues of track queries within a parallel decoder remains non-trivial.
@@
To thoroughly tackle these challenges, we propose a novel synergistic training strategy for the two tasks based on the tracking-by-attention paradigm, which we term SynCL. As presented in Fig. 1b, SynCL introduces a weight-shared parallel decoder during training, which removes the self-attention mechanism. Concretely, we propose a Task-specific Hybrid Matching module that additionally employs one-to-many assignment for object queries to uncover and preserve promising candidates overlooked by the self-attention mechanism in the parallel decoder. With more diverse and refined candidate representations from one-to-many label supervision, predictions from object queries are robust to over-deduplication. Additionally, we design a Dynamic Query Filtering module based on the Gaussian Mixture Model (GMM) that flexibly selects reliable object queries for the one-to-many assignment according to the prediction quality and optimization state of the model. Moreover, to get rid of constraints from self-centric attention on track queries and establish cross-task connections, we introduce Instance-aware Contrastive Learning that aligns joint object and track queries corresponding to identical ground truth targets while separating irrelevant queries in the latent space and thus provides synergistic learning with richer and quality features.
@@
In summary, our main contributions are as follows: (1) We reveal that the optimization difficulties of joint detection and tracking training lie in the imperceptible effect of the self-attention mechanism across different queries. (2) We propose SynCL, a synergistic training strategy compatible with any tracking-by-attention paradigm tracker to address multi-task learning challenges. SynCL implements dynamic filtering-based hybrid matching and instance-aware contrastive learning to improve the performance of both detection and tracking. (3) We empirically show that SynCL brings remarkable improvements over various tracking-by-attention baselines with acceptable extra training cost and establishes new state-of-the-art performance on the camera-based nuScenes MOT benchmark.
@@ 2 Related Work
Tracking-by-detection. Traditional tracking methods in 2D or 3D [2, 3, 13, 14] follow the Tracking-by-detection paradigm, which decouples detection and tracking, regarding tracking as a post-processing step. Early works associate targets with existing trajectories by handcrafted features such as appearance similarity [1], geometric distance [15, 16, 17, 18]. Tracking-by-detection paradigm achieves further advancements when combined with query-based detectors. Specifically, learning modules such as GNN [19, 20] and Transformers [21, 11] are employed to merge multifaceted information. However, detection and tracking are complementary yet non-independent tasks. This partitioned paradigm design does not effectively optimize both tasks simultaneously.
Multi-task learning in MOT. Building upon DETR [22] of encoding detection targets as object queries, MOTR [23] utilizes the track queries, enriched with prior features of object queries, to propagate across frames, enabling an end-to-end tracking framework. MUTR3D [4] and PF-Track [5] extend the end-to-end framework to multi-camera 3D MOT. However, sharing the same model parameters for detection and tracking within an end-to-end framework introduces optimization difficulties in multi-task learning. Subsequent works [9, 10, 11] attempt to address these difficulties. STAR-Track [24] introduces a Latent Motion Model (LMM) to account for changes in query appearance features across frames and ADA-Track [6] integrates the appearance and geometric clues of the track queries to refine object queries. Both of them increase inference speed. TQDTrack [25] introduces temporal query denoising to enhance the modeling of track queries. OneTrack [8] utilizes attention mask to block conflicting gradient flows, but its design is confined to [26]. Despite these efforts, a general strategy for optimization difficulties in joint detection and tracking training remains unsolved.
@@ 3 Methodology
Figure 3: Analysis of self-attention heatmap in the standard decoder. The annotation numbers of the heatmap are aligned with the ID numbers in Fig. 2.
@@ 3.2 Preliminaries
SynCL is designed to be integrated into any tracking-by-attention paradigm multi-view 3D tracker. These trackers typically consist of a convolutional network-based feature encoder (e.g., ResNet [27], VoVNet [28]), a transformer-based decoder, and a predictor head for object classes and boxes. The goal of 3D MOT is to generate detection boxes of the instances in the perceptual scene with consistent IDs. Given $N$ images $\mathbf{I}_{t}=\{\mathbf{I}_{t}^{i},i=1,2,\dots,N\}$ captured by surrounding cameras at timestamp $t$, the encoder first extracts image features $\mathbf{F}_{t}=\{\mathbf{F}_{t}^{i},i=1,2,\dots,N\}$ with 3D point position embeddings [29, 30]. A crucial aspect of realizing end-to-end tracking is to allow the prior latent encoding from the previous frame, track queries $\mathbf{Q}^{trk}_{t-1}$, to be updated iteratively by the decoder interacting with the current frame image features. To detect new-born targets, object queries $\mathbf{Q}^{obj}$ with a fixed number of $\mathbf{N}_{obj}$ are merged together and fed into the decoder:
@@
However, matching object queries to new-born objects contradicts the detection pre-training stage, substantially weakening the supervision signals for object queries. Consequently, when track queries are initialized with object queries from the previous frame, these low-quality priors inevitably exacerbate the optimization difficulties in inter-temporal tracking. We find that this suboptimal assignment stems from the inconspicuous nature of the self-attention mechanism, preventing synergistic training for both identity-agnostic detection and identity-aware tracking.
Figure 4: Overview of SynCL. SynCL is based on tracking-by-attention paradigm trackers, with two weight-shared parallel decoders: a S-decoder (standard decoder) and a C-decoder (devoid of self-attention layers). In C-decoder, hybrid matching with one-to-many and one-to-one assignment is applied for object queries and track queries, respectively. Besides, a dynamic filtering module is designed to flexibly select reliable object queries for the one-to-many assignment. With identical ground-truth matching, contrastive learning unifies the representations between object and track queries, co-facilitating multi-task learning for detection and tracking.
@@ 3.3 Task-specific Hybrid Matching
Parallel Decoders. We design two parallel decoders: a standard decoder (S-decoder) and a cross-attention-based decoder (C-decoder). A block in the standard decoder consists of a self-attention layer, a cross-attention layer and a feed-forward network (FFN). The embeddings of both object and track queries are concatenated and fed into the two decoders, denoted as $e^{l}_{s}$ and $e^{l}_{c}$, where $e^{l}=e^{l}_{obj}\cup e^{l}_{trk}$. The superscript $l=\{0,1,...,5\}$ indicates the index of the decoder block and we omit the notation of the frame index $t$ for simplicity. The initial inputs to both decoders are identical, i.e., $e^{0}_{s}=e^{0}_{c}$. In the $l_{th}$ block of the S-decoder, the state of both queries is first updated through the self-attention layer bidirectionally, which is formulated as:
$\hat{e}^{l}_{s}=e^{l}_{s}+{\rm Attention_{self}}(Query=Key=Value=e^{l}_{s})$ (3)
where $\rm Attention_{self}(\cdot)$ indicates the self-attention layer. Next, the two types of queries are interacted with multi-view image features, where object queries generate potential candidates through the positional priors, and track queries locate persistent targets utilizing representation prior:
$e^{l+1}_{s}={\rm FFN}(\hat{e}^{l}_{s}+{\rm Attention_{cross}}(Query=\hat{e}^{l}_{s},~{}Key=Value=\mathbf{F}))$ (4)
where $\rm Attention_{cross}(\cdot)$ indicates the cross-attention layer. The C-decoder retains the cross-attention layers and FFNs, sharing parameters with the S-decoder. In this setup, the cross-attention layers take image features as the sole source set (key and value), dynamically aggregating and updating both queries in the stacked decoder blocks:
$e^{l+1}_{c}={\rm FFN}(e^{l}_{c}+{\rm Attention_{cross}}(Query=e^{l}_{c},~{}Key=Value=\mathbf{F}))$ (5)
Due to the absence of mutual interaction between object and track queries, it is challenging for the C-decoder to remove duplicate candidates with proximal or occluded targets. Thus, we design distinct label assignment tailored to the characteristics of the two parallel decoders and the specific requirements of their respective tasks.
Hybrid Matching. End-to-end tracking-by-attention paradigm trackers rely on one-to-one assignment for both object and track queries. In the S-decoder, assignment of object queries $\mathbf{\sigma}^{obj}_{s}$ can be obtained by performing the Hungarian algorithm between predictions and new-born targets, while assignment of track queries $\mathbf{\sigma}^{trk}_{s}$ is consistent with the assigned IDs, which is formulated as:
@@
$\mathbf{\sigma}^{trk}_{s}$ =$=\left\{\sigma\mid\mathrm{ID}_{\sigma(i)}^{trk}=\mathrm{\hat{ID}}_{i},~{}\mathrm{\hat{ID}}\in g_{pst},i=1,2,\ldots,N\right\}$ (7)
where the $g_{new}$ and $g_{pst}$ are $M$ new-born and $N$ persistent ground truths (GTs). $\mathcal{C}_{match}$ is the matching cost between predictions and GTs. $\sigma(\cdot)$ is the optimal permutation of $M$ or $N$ indices, where $k_{th}$ ground truth target is assigned to $\sigma(k)_{th}$ prediction. Each query corresponds to a ground-truth target and vice versa. The track queries corresponding to disappearing targets are not assigned labels. Following [30, 31], we utilize a combination of the classification score and the 3D box as the matching cost metric, which is formulated as:
$C_{match}(b,p,g)={\rm FocalLoss}(p,\hat{p})+{\rm L1}(b,\hat{b})$ (8)
@@
where the $K$ is the maximum assignment constraint per GT for object queries. In order to obtain more positive sample pairs for contrastive learning, we do not match object queries with all ground-truth targets. In Sec. 4.3, we explore the impact of different matching variants in detail.
3.4 Dynamic Query Filtering
To effectively explore more high-quality candidate samples for the identity-agnostic matching in Hybrid Matching, we propose a cost-based filtering module with a dynamic threshold $\delta$, which filters out unreliable matching pairs whose cost exceeds $\delta$. Specifically, we use Gaussian Mixture Model (GMM) to unsupervisedly cluster matching sample pairs into reliable/unreliable clusters based on their matching cost $z$, which is formulated as:
$\bm{\pi},\bm{\mu},\bm{\Sigma}=\mathop{\arg\min}\sum_{z}-\log\sum_{i=1}^{N_{c}}\pi_{i}\mathcal{N}(z|\mu_{i},\Sigma_{i})$ (10)
where $\mathcal{N}(z|\mu_{i},\Sigma_{i})$ is the Gaussian probability density function with mean $\mu_{i}$ and covariance $\Sigma_{i}$, and $\pi_{i}$ is the weight for the $i$-th component, satisfying $\sum_{i=1}^{2}\pi_{i}=1$. The cluster with lower $\mu_{i}$ is considered as the reliable cluster and we use its mean $\mu_{reliable}=\min_{i}\mu_{i}$ as the dynamic threshold $\delta$. Notably, this dynamic design intelligently constructs a progressive learning schedule ranging from simple to complex depending on the optimization status of the model.